# AT 2025abao: the fourth luminous red nova in M 31

### v1

@@ 3.5 Bolometric light curve
The bolometric luminosity of AT 2025abao at the time of the -band maximum is , similar to the values reached by the comparison objects. However, its top luminosity () is reached about 67 d later, just before the fall from the plateau. This peak bolometric luminosity makes AT 2025abao the brightest LRN ever discovered in M 31 and, at the time of the peak, one of the most luminous objects within that galaxy. However, while the bolometric luminosity of AT 2025abao at peak is similar to that of the comparison objects and the fast rise resembles that of AT 2019zhd, the later evolution is completely different, as none of the LRNe considered presents a 70 d-lasting plateau. In this respect, AT 2025abao is more similar to AT 2018bwo (Blagorodnova et al., 2021; Pastorello et al., 2023), although the former is about three times fainter.
@@ 4 Spectral evolution
AT 2025abao was classified as a LRN on 2025 October 29, 10 days after the official discovery date and at a phase of d (Taguchi and Maeda, 2025b). However, earlier spectra from amateur astronomers are also available. The earliest, taken on 2025 October 27 (phase d), was presented by Skopal et al. (2025). One day later, another low-resolution spectrum was obtained by C. Balcon (phase d), and additional spectra were obtained with a nearly daily cadence (at phases , and d) by Taguchi and Maeda (2025b) and Skopal et al. (2025). These early spectra exhibit a blue continuum ( K) with superimposed narrow emission lines of the Balmer series, a feature typical of LRNe during the rise to maximum light (Pastorello et al., 2019a). The classification spectrum of Taguchi and Maeda (2025b), retrieved from the TNS, is shown in Fig. 9, where it is compared with spectra of LRNe AT 2019zhd and V1309 Sco at the maximum light to emphasize their overall similarity. Nonetheless, the spectral evolution of V1309 Sco is more rapid, and metal lines are already visible in absorption at these early epochs.
We conducted a spectroscopic follow-up of AT 2025abao, during which we collected 20 optical spectra spanning three months of evolution. The log of spectroscopic observations is given in Table 2, and the spectral time series is shown in Figure 10, left panel. The spectra from the 1.22m Galileo telescope equipped with a Boller & Chivens (B&C) spectrograph and the 1.82m Copernico telescope with an Echelle spectrograph were reduced with a custom-made, IRAF-based pipeline developed by A. Siviero, while the spectra from the 2.56m Nordic Optical Telescope (NOT) with ALFOSC, the 1.82m Copernico telescope with AFOSC, and the 10.4m Gran Telescopio Canarias (GTC) plus OSIRIS were reduced using the foscgui141414Foscgui is a graphic user interface aimed at extracting SN spectroscopy and photometry obtained with FOSC-like instruments. It was developed by E. Cappellaro. A package description can be found at http://sngroup.oapd.inaf.it/foscgui.html. pipeline. Finally, the spectra of the 3.58m Telescopio Nazionale Galileo (TNG) plus DOLORES spectra were hand-reduced using routine IRAF procedures.
@@
In the d spectrum of AT 2025abao, the first signs of molecular absorption bands: TiO features start to appear between 6000 and 8000 Å, in the form of two sawtooth drops in the continuum, sharp on the blue edge and then gradually rising towards the red edge. This trend of molecules formation is confirmed in the d spectrum. Molecular bands from VO and TiO are clearly present in the regions 6100-6700 Å and 7100-7800 Å, together with a rise at the start of the NIR, in the range 8400-8800 Å, also compatible with a molecular band. The continuum peaks at around 7800 Å (3700 K), with a temperature trend consistent with that expected for a moderately faint LRN. Balmer lines appear to be strong again in the d spectrum, while the continuum is now red and its temperature has decreased to 3600 K.
@@ 4.1 Spectral line identification
In the AT 2025abao spectrum, H shows a narrow P Cygni profile with a dominant absorption with a minimum blue-shifted by km s-1, and a weaker and narrow emission component. However, on the red side, a sharper counter P Cygni is also present. This feature is common in LRN spectra (see AT 2011kp; Pastorello et al. 2019a), and can be interpreted as outflowing material from the central system through the second Lagrangian point (Pejcha et al., 2016a) in a direction opposite to the observer. Other Balmer lines are detected (H, H, and H, and possibly also H), dominated by the emission component. Ca ii H&K are relatively broad (FWHM km s-1) and in pure absorption. Other notable absorption features are the Ca ii NIR triplet, and the Na i D doublet. All other features, while present and identified, are less strong.
In Fig. 13, we compare the full GTC+OSIRIS mid-resolution spectrum of AT 2025abao with that of AT 2011kp during late phases ( d) from Pastorello et al. (2019a), who also performed a similar line identification in a mid-resolution spectrum. The spectrum of AT 2025abao was taken 20 days after maximum, while that of AT 2011kp was obtained after more than three months. While the spectral continuum shows some differences, the most important metal lines are identified in both spectra. This is an indication of a much faster evolution of AT 2025abao. Thus, a similar set of lines is identified and visualized in both spectra: multiplets from Fe ii, Fe i, Mg i, Sc ii, Ba ii (the strongest line being 6497) and Ti ii are retrieved. Single lines from H, O i, Ca ii, and Na I are also found. In the red extreme, in a region not strongly contaminated by telluric absorption bands, we also identify the Na I 8183,8195 doublet, O i 7774 and 8446. Many lines, not marked in Fig. 13, can be attributed to a multitude of transitions from Ti ii, as suggested also by Pastorello et al. (2019a).
@@ 5 Discussion and conclusion
Usually, the progenitors of LRNe are found to be yellow stars, often supergiants in the cases of the most luminous events (Mauerhan et al., 2015; Blagorodnova et al., 2017; Pastorello et al., 2021c; Blagorodnova et al., 2021; Cai et al., 2022a). However, the association of AT 2025abao with WNTR23bzdiq, a slowly variable AGB star (K25), makes it an exception. K25 interpreted this object as a common-envelope event and predicted a merging event as the outcome. AGB stars are so inflated that the outer layers of the atmosphere can be considered like a hot vacuum. Therefore, the companion star - especially if the mass ratio is particularly large - could pass through it, remaining engulfed within the envelope and without significantly changing the global colour of the system, which would remain red, as observed by K25.
@@
K25 estimated a mass of the progenitor of based on evolutionary paths from MESA isochrones and stellar tracks. Using the empirical relation between absolute magnitude and the mass of the progenitor by Cai et al. (2022a) for an object with mag during the plateau phase171717Obviously, AT 2025abao does not exhibit a genuine secondary red peak, but instead a long plateau. However, the relation of Cai et al. (2022a) can still be used, considering the -band absolute magnitude at the end of the plateau. Consequently, in Appendix B, the Cai et al. (2022a) relation has been updated using also the AT 2025abao data., we obtain , in excellent agreement with their estimation. Based on the peak luminosity and similarities in the spectral evolution, we can also state that the progenitor mass of AT 2025abao must be similar to that of V838 Mon, which was (Pastorello et al., 2023; Tylenda, 2005) and slightly more massive than that of AT 2019zhd (6 for Pastorello et al. 2019b; 3.4 for Chen and Ivanova 2024).
@@ Acknowledgements.
AR acknowledges financial support from the GRAWITA Large Program Grant (PI D’Avanzo). AR, AP, GV, NER acknowledge financial support from the PRIN-INAF 2022 ”Shedding light on the nature of gap transients: from the observations to the models”. NER acknowledges support from the Spanish Ministerio de Ciencia e Innovación and the Agencia Estatal de Investigación 10.13039/501100011033 under the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. TMR acknowledges support from the Research Council of Finland project 350458 and the Cosmic Dawn Center, which is funded by the Danish National Research Foundation under grant DNRF140. MDS is funded by the Independent Research Fund Denmark (grant nr. 10.46540/2032-00022B). YZC is supported by the National Natural Science Foundation of China (NSFC, Grant No. 12303054), the National Key Research and Development Program of China (Grant No. 2024YFA1611603), the Yunnan Fundamental Research Projects (Grant Nos. 202401AU070063, 202501AS070078), and the International Centre of Supernovae, Yunnan Key Laboratory (No. 202302AN360001). Based on observations collected at Copernico and Schmidt telescopes (Asiago Mount Ekar, Italy) of the INAF – Osservatorio Astronomico di Padova, and at the Galileo telescope of the Padova Univesity (Asiago Mount Pennar, Italy). AR thanks Alessandro Fabris and Virginia Albanese for having participated to the observations with AFOSC. Support to ATLAS was provided by NASA grant NN12AR55G. Based on observations made with the NOT, owned and operated jointly by Aarhus University, the University of Turku, the University of Oslo, the University of Iceland and Stockholm University, under program 72-305 (PI Reguitti), and via the NUTS2 collaboration which is supported in part by the Instrument Centre for Danish Astrophysics, and the Finnish Centre for Astronomy with ESO via Academy of Finland grant nr. 306531. Based on observations made with the GTC. Based in part on observations made at the TNG, operated by INAF, under program A50TAC_41 (PI Valerin). All three telescopes are installed at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias, on the island of La Palma. We acknowledge the use of data from the Swift data archive. FR is grateful to iTelescope.Net for providing observing time to use their remote telescopes and to the AAVSO who made it possible to use the AAVSOnet telescopes.References
@@
- The luminous red nova AT 2018bwo in NGC 45 and its binary yellow supergiant progenitor. A&A 653, pp. A134. External Links: Document, 2102.05662, ADS entry Cited by: §1, §3.4, §3.5, §5, §5.
@@
- Forbidden hugs in pandemic times. III. Observations of the luminous red nova AT 2021biy in the nearby galaxy NGC 4631. A&A 667, pp. A4. External Links: Document, 2207.00734, ADS entry Cited by: Appendix B, §1, §5, §5, §5, footnote 11, footnote 17.
@@
- Light-curve Model for Luminous Red Novae and Inferences about the Ejecta of Stellar Mergers. ApJ 938 (1), pp. 5. External Links: Document, 2202.10478, ADS entry Cited by: §3.2.2, §5.
@@
- Forbidden hugs in pandemic times. I. Luminous red nova AT 2019zhd, a new merger in M 31. A&A 646, pp. A119. External Links: Document, 2011.10588, ADS entry Cited by: §3.2.1.
- Forbidden hugs in pandemic times. I. Luminous red nova AT 2019zhd, a new merger in M 31. A&A 646, pp. A119. External Links: Document, 2011.10588, ADS entry Cited by: §1, §2, §3.4.
- Luminous red novae: Stellar mergers or giant eruptions?. A&A 630, pp. A75. External Links: Document, 1906.00812, ADS entry Cited by: Figure 13, §4.1, §4.1, §4, footnote 11.
- Luminous red novae: Stellar mergers or giant eruptions?. A&A 630, pp. A75. External Links: Document, 1906.00812, ADS entry Cited by: §1, §3.4, §5, footnote 1.
@@
- Forbidden hugs in pandemic times. IV. Panchromatic evolution of three luminous red novae. A&A 671, pp. A158. External Links: Document, 2208.02782, ADS entry Cited by: §3.4, §3.5, §5, footnote 11.
@@
- The fate of the progenitors of Luminous Red Novae: Infrared detection of LRNe years after the outburst. A&A 706, pp. A154. External Links: Document, 2504.14592, ADS entry Cited by: §1, §5, §5, footnote 11.

### v2

@@ 3.5 Bolometric light curve
The bolometric luminosity of AT 2025abao at the time of the -band maximum is , similar to the values reached by the comparison objects. However, its top luminosity () is reached about 67 d later, just before the fall from the plateau. This peak bolometric luminosity makes AT 2025abao the brightest LRN ever discovered in M 31 and, at the time of the peak, one of the most luminous objects within that galaxy. However, while the bolometric luminosity of AT 2025abao at peak is similar to that of the comparison objects and the fast rise resembles that of AT 2019zhd, the later evolution is completely different, as none of the LRNe considered presents a 70 d-lasting plateau. In this respect, AT 2025abao is more similar to AT 2018bwo (Blagorodnova et al., 2021; Pastorello et al., 2023), although the former is about three times fainter. The total radiated energy, calculated by integrating the UV+optical+NIR bolometric curve between phases and d, is erg. This value has to be considered as a lower limit since we are missing the contributions from wavelengths longer than 2.5 m, and the energy emitted prior to phase d, i.e. the fast rise before the peak of the outburst. If we add the contribution from the filters from to between phases and d, and from the filters between and d, we obtain a total radiated energy of erg, which still has to be taken as an underestimation.
@@ 3.7 SPHEREx SED of the precursor
3.8 SPHEREx SED of the outburst
Between MJD 61027 and 61041 ( d to d), SPHEREx observed again the field of AT 2025abao, this time during the outburst stage. The corresponding IR SED, presented in Figure 9, shows an hot thermal continuum, with absorption bands at 0.82 m (visible also in the optical spectra, produced by VO and TiO molecules, as observed also in V838 Mon; Lynch et al. 2004), 0.94 m (due to TiO; Valenti et al., 1998), 1.1 m (from a combination of VO, CN and TiO molecules; Pastorello et al., 2023), 1.45 m (water vapours), and 2.4 m (from CO; Blagorodnova et al., 2021). We performed a Planckian fit to the reddening-corrected SED, which provides an excellent match up to the red tail with a BB having the following parameters: K, , and erg s-1. These parameters are consistent with those obtained from the optical filters at similar phases. No IR excess is detected up to 5 m, as a single BB is sufficient to fit the data even at the reddest wavelengths. Therefore, either dust is not present and had yet to be formed, or it is pre-existing and is much colder than 600 K, otherwise its emission would have been visible as an IR excess at such long wavelengths.
@@ 4 Spectral evolution
AT 2025abao was classified as a LRN on 2025 October 29, 10 days after the official discovery date and at a phase of d (Taguchi and Maeda, 2025b). However, earlier spectra from amateur astronomers are also available. The earliest, taken on 2025 October 27 (phase d), was presented by Skopal et al. (2025). One day later, another low-resolution spectrum was obtained by C. Balcon (phase d), and additional spectra were obtained with a nearly daily cadence (at phases , and d) by Taguchi and Maeda (2025b) and Skopal et al. (2025). These early spectra exhibit a blue continuum ( K) with superimposed narrow emission lines of the Balmer series, a feature typical of LRNe during the rise to maximum light (Pastorello et al., 2019). The classification spectrum of Taguchi and Maeda (2025b), retrieved from the TNS, is shown in Fig. 10, where it is compared with spectra of LRNe AT 2019zhd and V1309 Sco at the maximum light to emphasize their overall similarity. Nonetheless, the spectral evolution of V1309 Sco is more rapid, and metal lines are already visible in absorption at these early epochs.
We conducted a spectroscopic follow-up of AT 2025abao, during which we collected 21 optical spectra spanning more than three months of evolution. The log of spectroscopic observations is given in Table 2, and the spectral time series is shown in Figure 11, left panel. The spectra from the 1.22m Galileo telescope equipped with a Boller & Chivens (B&C) spectrograph and the 1.82m Copernico telescope with an Echelle spectrograph were reduced with a custom-made, IRAF-based pipeline developed by A. Siviero, while the spectra from the 2.56m Nordic Optical Telescope (NOT) with ALFOSC, the 1.82m Copernico telescope with AFOSC, and the 10.4m Gran Telescopio Canarias (GTC) plus OSIRIS were reduced using the foscgui141414Foscgui is a graphic user interface aimed at extracting SN spectroscopy and photometry obtained with FOSC-like instruments. It was developed by E. Cappellaro. A package description can be found at http://sngroup.oapd.inaf.it/foscgui.html. pipeline. Finally, the spectra of the 3.58m Telescopio Nazionale Galileo (TNG) plus DOLORES spectra were hand-reduced using routine IRAF procedures.
@@
In the d spectrum of AT 2025abao, the first signs of molecular absorption bands: TiO features start to appear between 6000 and 8000 Å, in the form of two sawtooth drops in the continuum, sharp on the blue edge and then gradually rising towards the red edge. This trend of molecules formation is confirmed in the d spectrum. Molecular bands from VO and TiO are clearly present in the regions 6100-6700 Å and 7100-7800 Å, together with a rise at the start of the NIR, in the range 8400-8800 Å, also compatible with a molecular band. The continuum peaks at around 7800 Å (3700 K), with a temperature trend consistent with that expected for a moderately faint LRN. Balmer lines appear to be strong again in the and d spectra, while the continuum is now red and its temperature has decreased to 3600 and 3500 K, respectively.
@@ 4.1 Spectral line identification
In the AT 2025abao spectrum, H shows a narrow P Cygni profile with a dominant absorption with a minimum blue-shifted by km s-1, and a weaker and narrow emission component. However, on the red side, a sharper counter P Cygni is also present. This feature is common in LRN spectra (see AT 2011kp; Pastorello et al. 2019), and can be interpreted as outflowing material from the central system through the second Lagrangian point (Pejcha et al., 2016a) in a direction opposite to the observer. Other Balmer lines are detected (H, H, and H, and possibly also H), dominated by the emission component. Ca ii H&K are relatively broad (FWHM km s-1) and in pure absorption. Other notable absorption features are the Ca ii NIR triplet, and the Na i D doublet. All other features, while present and identified, are less strong.
In Fig. 14, we compare the full GTC+OSIRIS mid-resolution spectrum of AT 2025abao with that of AT 2011kp during late phases ( d) from Pastorello et al. (2019), who also performed a similar line identification in a mid-resolution spectrum. The spectrum of AT 2025abao was taken 20 days after maximum, while that of AT 2011kp was obtained after more than three months. While the spectral continuum shows some differences, the most important metal lines are identified in both spectra. This is an indication of a much faster evolution of AT 2025abao. Thus, a similar set of lines is identified and visualized in both spectra: multiplets from Fe ii, Fe i, Mg i, Sc ii, Ba ii (the strongest line being 6497) and Ti ii are retrieved. Single lines from H, O i, Ca ii, and Na I are also found. In the red extreme, in a region not strongly contaminated by telluric absorption bands, we also identify the Na I 8183,8195 doublet, O i 7774 and 8446. Many lines, not marked in Fig. 14, can be attributed to a multitude of transitions from Ti ii, as suggested also by Pastorello et al. (2019).
@@ 5 Discussion and conclusion
Usually, the progenitors of LRNe are found to be yellow stars, often supergiants in the cases of the most luminous events (Mauerhan et al., 2015; Blagorodnova et al., 2017; Pastorello et al., 2021b; Blagorodnova et al., 2021; Cai et al., 2022a). However, the association of AT 2025abao with WNTR23bzdiq, a slowly variable AGB star (K25), makes it an exception. K25 interpreted this object as a common-envelope event and predicted a merging event as the outcome. AGB stars are so inflated that the outer layers of the atmosphere can be considered like a hot vacuum. Therefore, the companion star - especially if the mass ratio is particularly large - could pass through it, remaining engulfed within the envelope and without significantly changing the global colour of the system, which would remain red, as observed by K25.
@@
K25 estimated a mass of the progenitor of based on evolutionary paths from MESA isochrones and stellar tracks. Using the empirical relation between absolute magnitude and the mass of the progenitor by Cai et al. (2022a) for an object with mag during the plateau phase171717Obviously, AT 2025abao does not exhibit a genuine secondary red peak, but instead a long plateau. However, the relation of Cai et al. (2022a) can still be used, considering the -band absolute magnitude at the end of the plateau. Consequently, in Appendix B, the Cai et al. (2022a) relation has been updated using also the AT 2025abao data., we obtain , in excellent agreement with their estimation. Based on the peak luminosity and similarities in the spectral evolution, we can also state that the progenitor mass of AT 2025abao must be similar to that of V838 Mon, which was (Pastorello et al., 2023; Tylenda, 2005) and slightly more massive than that of AT 2019zhd (3.4 for Pastorello et al. 2021a; 6 for Chen and Ivanova 2024).
To provide more physical parameters, we considered the analytical expressions derived by Matsumoto and Metzger (2022). In particular, their Equation 16 allows to estimate the mass of the ejected material by measuring the duration of the plateau and the bulk velocity of the ejecta . For the plateau length, we assume days, as in the -band it started one week before maximum light, and ended at d. For the ejecta velocity, we adopted the FWHM of the Hlines during the plateau (as in Cai et al. 2022a), precisely in the and d spectra, which give km s-1. Under these assumptions, we calculate and ejecta mass of 0.63 .
@@ Acknowledgements.
AR acknowledges financial support from the GRAWITA Large Program Grant (PI D’Avanzo). AR, AP, GV, NER acknowledge financial support from the PRIN-INAF 2022 ”Shedding light on the nature of gap transients: from the observations to the models”. NER acknowledges support from the Spanish Ministerio de Ciencia e Innovación and the Agencia Estatal de Investigación 10.13039/501100011033 under the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. TMR acknowledges support from the Research Council of Finland project 350458 and the Cosmic Dawn Center, which is funded by the Danish National Research Foundation under grant DNRF140. MDS is funded by the Independent Research Fund Denmark (grant nr. 10.46540/2032-00022B). YZC is supported by the National Natural Science Foundation of China (NSFC, Grant No. 12303054), the National Key Research and Development Program of China (Grant No. 2024YFA1611603), the Yunnan Fundamental Research Projects (Grant Nos. 202401AU070063, 202501AS070078), and the International Centre of Supernovae, Yunnan Key Laboratory (No. 202302AN360001). Based on observations collected at Copernico and Schmidt telescopes (Asiago Mount Ekar, Italy) of the INAF – Osservatorio Astronomico di Padova, and at the Galileo telescope of the Padova Univesity (Asiago Mount Pennar, Italy). AR thanks Alessandro Fabris and Virginia Albanese for having participated to the observations with AFOSC. Support to ATLAS was provided by NASA grant NN12AR55G. Based on observations made with the NOT, owned and operated jointly by Aarhus University, the University of Turku, the University of Oslo, the University of Iceland and Stockholm University, under program 72-305 (PI Reguitti), and via the NUTS2 collaboration which is supported in part by the Instrument Centre for Danish Astrophysics, and the Finnish Centre for Astronomy with ESO via Academy of Finland grant nr. 306531. Based on observations made with the GTC. Based in part on observations made at the TNG, operated by INAF, under program A50TAC_41 (PI Valerin). All three telescopes are installed at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias, on the island of La Palma. We acknowledge the use of data from the Swift data archive. FR is grateful to iTelescope.Net for providing observing time to use their remote telescopes and to the AAVSO who made it possible to use the AAVSOnet telescopes. SPHEREx is a joint project of the Jet Propulsion Laboratory and the California Institute of Technology, and is funded by the National Aeronautics and Space Administration.References
@@
- Evidence for the merger hypothesis in V4332 Sgr: a low 12C/13C ratio and multiple outbursts. arXiv e-prints, pp. arXiv:2602.14781. External Links: 2602.14781, ADS entry Cited by: §3.2.2.
@@
- The luminous red nova AT 2018bwo in NGC 45 and its binary yellow supergiant progenitor. A&A 653, pp. A134. External Links: Document, 2102.05662, ADS entry Cited by: §1, §3.4, §3.5, §3.8, §5, §5.
@@
- Forbidden hugs in pandemic times. III. Observations of the luminous red nova AT 2021biy in the nearby galaxy NGC 4631. A&A 667, pp. A4. External Links: Document, 2207.00734, ADS entry Cited by: Appendix B, §1, §5, §5, §5, §5, footnote 11, footnote 17.
@@
- 0.8-13 Micron Spectroscopy of V838 Monocerotis and a Model for Its Emission. ApJ 607 (1), pp. 460–473. External Links: Document, ADS entry Cited by: §3.8.
@@
- Light-curve Model for Luminous Red Novae and Inferences about the Ejecta of Stellar Mergers. ApJ 938 (1), pp. 5. External Links: Document, 2202.10478, ADS entry Cited by: §3.2.2, §5, §5.
@@
- Forbidden hugs in pandemic times. I. Luminous red nova AT 2019zhd, a new merger in M 31. A&A 646, pp. A119. External Links: Document, 2011.10588, ADS entry Cited by: §1, §2, §3.2.1, §3.4, §5.
- Luminous red novae: Stellar mergers or giant eruptions?. A&A 630, pp. A75. External Links: Document, 1906.00812, ADS entry Cited by: §1, §3.4, Figure 14, §4.1, §4.1, §4, footnote 1, footnote 11.
@@
- Forbidden hugs in pandemic times. IV. Panchromatic evolution of three luminous red novae. A&A 671, pp. A158. External Links: Document, 2208.02782, ADS entry Cited by: §3.4, §3.5, §3.8, §5, footnote 11.
@@
- The fate of the progenitors of luminous red novae: Infrared detection of LRNe years after the outburst. A&A 706, pp. A154. External Links: Document, ADS entry Cited by: §1, §5, §5, footnote 11.
@@
- Spectral Synthesis of TiO Lines. ApJ 498 (2), pp. 851–862. External Links: Document, ADS entry Cited by: §3.8.